OpenStudy (anonymous):

karla has spent $39 on food she bought milk cheese and meat the milk cost half as much as the cheese and the meat cost four times as much as the cheese. which equation would help you find the cost of the cheese?

OpenStudy (hannaha):

what's the question?

OpenStudy (anonymous):

I guess they want an equation out of that problem

OpenStudy (igreen):

milk = 1/2 of cheese meat = 4 times cheese You could write something like: c + 1/2c + 4c = 39

OpenStudy (anonymous):

thank you!!!